# Comparing Variable Selection and Model Averaging Methods for Logistic Regression

### v1

@@ Discussion
Our study is a conceptual replication of the work of Porwal and Raftery [PorwalRaftery_2022], who conducted perhaps the most extensive comparison to date of methods for variable selection under model uncertainty, focusing on linear regression. Their simulations, which were also based directly on real datasets, showed that adaptive Bayesian model averaging (BMA) methods using $g$-priors performed best across estimation, inference, and prediction tasks. The $g=\sqrt{n}$ and empirical Bayes local (EB-local) priors performed particularly well. Our design closely follows theirs but extends it to logistic regression, where separation and nonlinearity present additional challenges. The strong performance of $g$-prior-based and EB-local BMA methods in our study suggests that the advantages of adaptive priors extend beyond linear models. However, since separation does not arise in linear models, our results also reveal an additional strength of the EB-local method: it remains robust across models and data conditions, including those affected by separation.
@@
Many comparisons of variable selection methods have been made for linear regression, but few have been made specifically for logistic regression, where issues such as separation and nonlinearity introduce additional challenges. One of the most extensive earlier comparisons was carried out by Li and Clyde [li_clyde_2018], who evaluated a broad class of $g$-priors using simulated datasets with correlated and uncorrelated predictors and varying levels of sparsity in the generating model. The priors they examined included hyper-$g$, hyper-$g/n$, beta-prime, CCH, robust, intrinsic, benchmark, and EB-local priors. All of these priors were also part of our analysis. Li and Clyde found that no single prior performed best across all criteria; some variants performed better for model selection, while others performed better for prediction. However, their simulations were based on synthetic designs and did not include separation or comparisons with non-$g$-prior methods. Our study complements theirs by evaluating these priors alongside additional Bayesian and penalized likelihood methods in empirically grounded simulations. Consistent with their results, we found that adaptive and mixture-based $g$-priors perform well under no separation. However, we also observed a marked decline in their performance under separation, a scenario Li and Clyde did not consider. In this setting, the EB-local prior maintained strong performance across metrics, further supporting its value as a versatile default for logistic regression.
@@
Two smaller studies also compared methods for logistic regression and provide useful points of reference. Wang, Zhang, and Bakhai [wang2004comparison] compared BMA with classical stepwise procedures in two small simulation settings ($n=500$, $p=25$). In both settings, one with no true predictors and one with two predictors, BMA identified the correct model consistently, while stepwise procedures tended to select spurious predictors. Our results are fully consistent with theirs. BMA methods outperformed stepwise selection across all conditions. Wu, Ferreira, and Gompper [WuEtAl2016] compared the fixed $g=4$ prior, the hyper-$g/n$ prior, and BIC in simulations with eight predictors and varying sample sizes. Overall, the three methods performed similarly. The $g$-prior variants performed slightly better when predictors were present in the data-generating model, while BIC performed better for the empty data-generating model. Our results for the inference metric (AUPRC) are similar to theirs, but we also found that the fixed $g=4$ prior performed worse than the hyper-$g/n$ and BIC approaches in terms of estimation metrics (RMSE and MIS) in the non-separation scenario. This suggests that estimation accuracy is an important dimension of performance for $g$-priors, particularly under separation.
@@
Indeed, the sharp decline in the overall performance of $g$-prior-based methods for datasets showing separation is primarily driven by reduced point and interval estimation accuracy, as indicated by higher root mean squared errors and mean interval scores. Nevertheless, inference and predictive performance remain strong, with relatively stable areas under the precision–recall curve and Brier scores. This pattern aligns with the theoretical results of Li and Clyde [li_clyde_2018], who showed that posterior distributions under $g$-priors may not be well-defined when maximum likelihood estimates do not exist (as in the case of separation). However, model-averaged inference and predictive quantities remain valid.
@@
Several other Bayesian variable selection methods for logistic regression have recently been proposed. However, we did not include them in our comparison because we could not find any publicly available software implementations. These methods include the power-expected-posterior prior of Fouskakis, Ntzoufras, and Perrakis [Fouskakis2018]; the nonlocal prior approach of Shi, Lim, and Maiti [Shi2019]; the variational Bayes method of Zhang, Xu, and Zhang [Zhang2019]; and the environmental DNA–motivated model of Griffin et al. [Griffin2020]. Earlier work by Liao and Chin [LiaoChin2007] proposed a parametric bootstrap method tailored to high-dimensional ($p\gg n$) settings. This method was implemented in the GeneLogit package, which is no longer available. Once accessible implementations exist, evaluating these and other emerging methods would further strengthen comparative assessments of Bayesian variable selection for logistic regression.
@@
Our analysis also focused narrowly on logistic regression methods. We did not compare logistic regression with other machine learning algorithms because doing so would have expanded the scope of the study considerably. Song et al. [Song2021], who reviewed 24 studies comparing logistic regression with various machine learning methods for predicting acute kidney injury, found broadly similar performance between the two approaches, though there was greater variability among machine learning methods. Additionally, Sur and Candès [SurCandes2019] showed that standard large-sample theory for significance testing in logistic regression breaks down when the number of variables is large, even if it remains smaller than the sample size. They proposed an alternative asymptotic framework. It would be valuable to examine how the methods evaluated here behave under those high-dimensional conditions.
@@
Finally, our comparison addressed model uncertainty arising from uncertainty about which variables to include, but not other forms of model uncertainty such as nonlinear effects, variable interactions, or alternative link functions like the complementary log-log function. Extending the present framework to these settings would provide a more complete picture of model uncertainty in generalized linear models. Likewise, future work could explore the impact of different model space priors. In this study, we used a truncated beta-binomial(1,1) prior that assigns equal probability across model sizes up to n2n-2. However, alternative priors that more strongly penalize model complexity or encode prior information about sparsity could lead to different conclusions.
@@
In summary, our study provides a comprehensive, preregistered, empirical comparison of 28 established methods for variable selection and inference under model uncertainty in logistic regression. The EB-local prior was the most robust method across separation and non-separation conditions, offering a balance of accuracy, stability, and computational efficiency. Penalized methods, such as SCAD and MCP, are strong alternatives when separation is likely and quantifying model uncertainty is deemed unnecessary. Together, these findings and the preregistered, data-based design of the study demonstrate the feasibility of transparent, reproducible simulation studies in statistical methodology.
@@
Table 1 provides an overview of the included methods. By methods, we also refer to different implementation options within R packages. For instance, the BAS package offers several parameter priors that have been shown to yield varying performance [cf. li_clyde_2018, PorwalRaftery_2022]; we therefore treat each of these priors as a distinct method.
@@
Method Authors Implementation Function
Hyperg-g [method_BAS] BAS-V1.7.5 bas.glm(..., betaprior = hyper.g())
@@
$g=\sqrt{n}$ [method_BAS] BAS-V1.7.5 bas.glm(..., betaprior = g.prior(sqrt(n)))
$g=\max(n,\,p^{2})$ [method_BAS] BAS-V1.7.5 bas.glm(..., betaprior = g.prior(max(n, p^2)))
$g=4$ [method_BAS] BAS-V1.7.5 bas.glm(..., betaprior = g.prior(4))
CCH [method_BAS] BAS-V1.7.5 bas.glm(..., betaprior = CCH(1, n, 0))
Robust [bayarri2012criteria] BAS-V1.7.5 bas.glm(..., betaprior = robust())
Intrinsic [method_BAS] BAS-V1.7.5 bas.glm(..., betaprior = intrinsic())
Beta-prime [method_BAS] BAS-V1.7.5 bas.glm(..., betaprior = beta.prime())
EB-local [method_BAS] BAS-V1.7.5 bas.glm(..., betaprior = EB.local())
EB-global [method_BAS] BAS-V1.7.5 bas.glm(..., betaprior = EB.global())
AIC [method_BAS] BAS-V1.7.5 bas.glm(..., betaprior = aic.prior())
BIC [method_BAS] BAS-V1.7.5 bas.glm(..., betaprior = bic.prior())
Ridge [method_glmnet] glmnet-V4.1.9 cv.glmnet(..., alpha = 0)
Elastic net [method_glmnet] glmnet-V4.1.9 cv.glmnet(..., alpha = 1/2)
LASSO [method_glmnet] glmnet-V4.1.9 cv.glmnet(..., alpha = 1)
$p$-value [method_glm] stats-V4.5.0 glm(...)
$p<.05$ [method_glm] stats-V4.5.0 glm(...)
$p<.005$ [method_glm] stats-V4.5.0 glm(...)
Forward [method_glm] stats-V4.5.0 step(glm(...), direction = "forward"
Backward [method_glm] stats-V4.5.0 step(glm(...), direction = "backward"
Both [method_glm] stats-V4.5.0 step(glm(...), direction = "both"
Spike and Slab [tuchler2008bayesian] BoomSpikeSlab-V1.2.6 logit.spike(..., prior = LogitZellnerPrior(...))
MCP [method_ncvreg] ncvreg-V3.15.0 cv.ncvreg(..., penalty = "MCP")
SCAD [method_ncvreg] ncvreg-V3.15.0 cv.ncvreg(..., penalty = "SCAD")
Induced Smoothed lasso [method_islasso] islasso-V1.5.2 islasso(...)
Firth’s Bias-Reduced [method_logistf] logistf-V1.26.1 logistf(...)
BMA [method_BMA] BMA-V3.18.20 bic.glm(...)
For the sake of computational efficiency, we limited the number of posterior samples for BAS and BoomSpikeSlab to 10,000, matching the number of samples used in earlier large scale simulation studies [e.g., PorwalRaftery_2022, porwal2022_modelpriors]. In addition, for BAS we did not vary the model space prior. Instead, we fixed the model space prior to a truncated beta-binomial($\alpha=1,\,\beta=1$) so that models with n2n-2 predictors or more received a prior probability of 0. For datasets with $n>p$, this reduces to an untruncated beta binomial prior that is uniform over the model size. In the preregistration, we indicated that we would conduct initial simulation studies to determine which model-space priors to use. However, due to computational considerations, we ultimately decided to retain the default choices described above.
@@
We began by screening online data repositories (e.g., Kaggle, Open Psychometrics, University of California Irvine Machine Learning Repository), datasets available in software packages (e.g., R, the JASP data library JASP2025), educational materials [e.g., hosmer2013applied, agresti2010analysis], and additional resources such as published research articles for suitable empirical datasets. This initial screening stage yielded 51 candidate datasets. From these, we excluded those that were not readily accessible (i.e., required substantial preprocessing or assumptions about the variables) and those that were unsuitable for logistic regression (e.g., highly imbalanced outcome variables or clear assumption violations). This selection process resulted in 24 candidate datasets for the simulation. These candidate data sets correspond to those specified in the preregistration document.
In addition to ensuring diversity in data dimensions and subject areas, we also assessed the suitability of each dataset for the data-generating process (DGP; described in the following subsection), as we anticipated that some might fail to produce a valid data-generating model (DGM). Each candidate dataset was evaluated using a DGP function. A dataset was automatically included if it (a) produced no computational issues or warnings, (b) exhibited no separation problems (thus requiring no special handling; see the following subsection), (c) yielded a DGM that was a proper subset of the full predictor set, (d) achieved a pseudo-$R^{2}$ greater than $0.1$, and (e) was not too thematically similar to another dataset already included. Based on these criteria, 11 datasets were retained for inclusion, while 8 were excluded for failing one or more conditions.
@@
burn 388 6 [hosmer2013applied, R-aplore3]
telescope 19,020 10 [mitgandhi2023logistic]
dyslexia 118 11 [Dbska2021, osfDbska2021]
housing 545 11 [yasserh2023housing]
heart 1,025 13 [johnsmith88_heart_disease]
covid19 468 25 [osf-covid19]
voting 345 34 [Brandt-2021]
divorce 170 54 [UCIdivorce]
nki70 144 70 [nki70-R-penalized]
musk 6,598 166 [chapman1994musk]
Singh 102 6,033 [singh2002]
@@
      • ISIS: A method that iteratively selects important variables by ranking them according to their marginal correlations with the outcome [fan2010sure]. It then performs conditional screening given previously selected variables.
@@
    Simulation process: After obtaining the coefficients for the selected predictors, we calculate the success probability for each observation in the dataset, that is $\pi_{i}=1/(1+\exp{\left(-\bm{\hat{\beta}}^{\intercal}\mathbf{x}_{i}\right)})$, with $i=1,\ldots,n$, where we bound the linear predictor $\bm{\beta}^{\intercal}\mathbf{x}_{i}$ between 10-10 and $10$ to avoid numerical instability. Finally, we generate $M$ random binary outcomes for each observation, that is, $y_{i}^{(m)}\sim\text{Bernoulli}(\pi_{i})$, for $m=1,\ldots,M$ and $i=1,\ldots,n$.
Simulation Design & Metrics
@@
We simulated 100 binary outcome variables for each empirical dataset (Table 2) by drawing binary outcome vectors $Y_{i}$ from the corresponding data-generating model (see previous subsection), while keeping the predictor matrix $\bm{X}$ fixed. This procedure yielded a total of $11\times 100$ simulated datasets. Each dataset was analyzed using all 28 candidate methods (Table 1). To ensure robustness, we applied 5-fold cross-validation when analyzing each simulated dataset with each method.
We monitored whether separation occurred in any dataset at any fold of the cross-validation using the R package detectseparation [detectseparation_package]. Out of the 1,100 simulated datasets, 463 (42%) exhibited separation. For the covid19 dataset, 34 out of 100 simulated datasets showed separation; for heart, 28 out of 100; and for housing, 1 out of 100. All simulated datasets showed separation for the divorce, nki70, voting, and Singh datasets, whereas none exhibited separation for the burn, dyslexia, musk, and telescope datasets.
@@
A complete absence of results (i.e., total failure) for an entire method occurred only in cases where the data exhibited separation. For the Singh dataset, all replications failed for BMA, the Induced Smoothed Lasso, Firth’s Bias-Reduced method, and all stepwise and $p$-value–based methods (nine methods in total). For the Voting dataset, all replications failed for BMA, the Induced Smoothed Lasso, Firth’s Bias-Reduced method, forward selection, and all $p$-value–based methods (seven methods in total). For the divorce and nki70 datasets, the $p<0.05$ and $p<0.005$ methods failed completely (two methods each). Finally, for the covid19 dataset, the $p<0.05$ method failed. Overall, these failures accounted for 7.6% of the entire simulation and 19.2% of all instances involving separation.
The performance measures used to compare the methods are described below. Each measure was averaged across the 100 repetitions for each simulation based on the 11 empirical datasets. The results presented in Figs. 1 and 2 represent averages across these 11 simulations. In addition, we report two supplementary metrics that, while not used for ranking the methods, are important for interpreting the results.
For the presentation of the main results in Figs. 1 and 2, all performance metrics were first averaged over repetitions for each dataset–method pair, then standardized as log-ratios relative to a fixed reference method (the Spike–and–Slab). These standardized values were exponentiated for display, such that a score of 1 denotes parity with the reference method, values below 1 indicate better performance, and values above 1 indicate worse performance relative to the reference method. The overall Full Score corresponds to the geometric mean across all performance metrics. The Available Score is computed on the subset of metrics provided by each method, and the Partial Score is restricted to the core metrics shared across all methods (RMSE and prediction Brier score). The methods shown in Figs. 1 and 2 are ranked according to the Partial Score. The additional metrics were not standardized.
@@
  • Interval Score (IS): The interval score [Gneiting2007] is a measure to assess the quality of confidence/credible intervals for the coefficients. It is calculated as the interval width penalized by the missed coverage when the true value lies outside the interval. The smaller interval score the better.
@@
    where $l$ and $u$ are respectively the lower and upper bound of the confidence/credible interval, and $2/\alpha$ is a penalization coefficient for the missed coverage, with $\alpha=0.05$.
@@
  • Brier score: The Brier score [brier1950verification] is a measure used to quantify the accuracy of binary probabilistic predictions on held-out test data. It is defined as the mean squared difference between predicted probabilities and the actual binary outcomes. Lower Brier scores indicate better calibrated and more accurate predictions:
@@
  • CPU time: The time in seconds that is taken to run a simulation on one dataset using a specific method over 100 repetitions. The values presented in Figs. 1 and 2 are in minutes.
  • The proportion of failed models (i.e., analyses that did not produce valid results).

### v2

@@ Discussion
Our study is a conceptual replication of the work of Porwal and Raftery [81], who conducted perhaps the most extensive comparison to date of methods for variable selection under model uncertainty, focusing on linear regression. Their simulations, which were also based directly on real datasets, showed that adaptive Bayesian model averaging (BMA) methods using $g$-priors performed best across estimation, inference, and prediction tasks. By adaptive BMA methods, we mean BMA methods in which the parameter priors and/or the model space prior are themselves estimated from the data. More specifically here, when $g$-priors are used, it refers to methods in which $g$ is estimated from or otherwise adapts to the data. This should not be confused with the adaptive MCMC algorithms sometimes used in Bayesian statistics.
The $g=\sqrt{n}$ and empirical Bayes local (EB-local) priors performed particularly well in that study [81]. Our design closely follows theirs but extends it to logistic regression, where separation and nonlinearity present additional challenges. The strong performance of $g$-prior-based and EB-local BMA methods in our study suggests that the advantages of adaptive priors extend beyond linear models. However, since separation does not arise in linear models, our results also reveal an additional strength of the EB-local method: it remains robust across models and data conditions, including those affected by separation.
@@
Many comparisons of variable selection methods have been made for linear regression, but few have been made specifically for logistic regression, where issues such as separation and nonlinearity introduce additional challenges. One of the most extensive earlier comparisons was carried out by Li and Clyde [71], who evaluated a broad class of $g$-priors using simulated datasets with correlated and uncorrelated predictors and varying levels of sparsity in the generating model. The priors they examined included hyper-$g$, hyper-$g/n$, beta-prime, CCH, robust, intrinsic, benchmark, and EB-local priors. All of these priors were also part of our analysis. Li and Clyde found that no single prior performed best across all criteria; some variants performed better for model selection, while others performed better for prediction. However, their simulations were based on synthetic designs and did not include separation or comparisons with non-$g$-prior methods. Our study complements theirs by evaluating these priors alongside additional Bayesian and penalized likelihood methods in empirically grounded simulations. Consistent with their results, we found that adaptive and mixture-based $g$-priors perform well under no separation. However, we also observed a marked decline in their performance under separation, a scenario Li and Clyde did not consider. In this setting, the EB-local prior maintained strong performance across metrics, further supporting its value as a versatile default for logistic regression.
@@
Two smaller studies also compared methods for logistic regression and provide useful points of reference. Wang, Zhang, and Bakhai [105] compared BMA with classical stepwise procedures in two small simulation settings ($n=500$, $p=25$). In both settings, one with no true predictors and one with two predictors, BMA identified the correct model consistently, while stepwise procedures tended to select spurious predictors. Our results are fully consistent with theirs. BMA methods outperformed stepwise selection across all conditions. Wu, Ferreira, and Gompper [108] compared the fixed $g=4$ prior, the hyper-$g/n$ prior, and BIC in simulations with eight predictors and varying sample sizes. Overall, the three methods performed similarly. The $g$-prior variants performed slightly better when predictors were present in the data-generating model, while BIC performed better for the empty data-generating model. Our results for the inference metric (AUPRC) are similar to theirs, but we also found that the fixed $g=4$ prior performed worse than the hyper-$g/n$ and BIC.BAS approaches in terms of estimation metrics (RMSE and MIS) in the non-separation scenario. This suggests that estimation accuracy is an important dimension of performance for $g$-priors, particularly under separation.
@@
Indeed, the sharp decline in the overall performance of $g$-prior-based methods for datasets showing separation is primarily driven by reduced point and interval estimation accuracy, as indicated by higher root mean squared errors and mean interval scores. Nevertheless, inference and predictive performance remain strong, with relatively stable areas under the precision–recall curve and Brier scores. This pattern aligns with the theoretical results of Li and Clyde [71], who showed that posterior distributions under $g$-priors may not be well-defined when maximum likelihood estimates do not exist (as in the case of separation). However, model-averaged inference and predictive quantities remain valid.
@@
Several other Bayesian variable selection methods for logistic regression have recently been proposed. However, we did not include them in our comparison because we could not find any publicly available software implementations. These methods include the power-expected-posterior prior of Fouskakis, Ntzoufras, and Perrakis [34]; the nonlocal prior approach of Shi, Lim, and Maiti [93]; the variational Bayes method of Zhang, Xu, and Zhang [112]; and the environmental DNA–motivated model of Griffin et al. [47]. For similar reasons, we did not include the test-based Bayes factor methods for probit regression proposed by Hu and Johnson [57] and logistic regression by Held, Sabanès Bovè, and Gravestock [51]. Earlier work by Liao and Chin [73] proposed a parametric bootstrap method tailored to high-dimensional ($p\gg n$) settings. This method was implemented in the GeneLogit package, which is no longer available. Similarly, methods using the Laplace Power Exponential Prior (LPEP) were shown in simulations to be competitive [83], although an R package on CRAN is not available at this time. Once accessible implementations exist, evaluating these and other emerging methods would further strengthen comparative assessments of Bayesian variable selection for logistic regression.
Most of the Bayesian methods we have considered involve model averaging, while most of the frequentist methods focus on model selection. This reflects the historical development of the area. However, in the past decade there has been increasing attention paid to frequentist model averaging (FMA) [52, 21, 96]. FMA itself involves many choices and hence has many variants, and the associated publicly available software is less developed than for the methods we have focused on here. However, it would be of interest to extend the present study to include FMA methods.
@@
Our analysis also focused narrowly on logistic regression methods. We did not compare logistic regression with other machine learning algorithms because doing so would have expanded the scope of the study considerably. Song et al. [95], who reviewed 24 studies comparing logistic regression with various machine learning methods for predicting acute kidney injury, found broadly similar performance between the two approaches, though there was greater variability among machine learning methods. Additionally, Sur and Candès [99] showed that standard large-sample theory for significance testing in logistic regression breaks down when the number of variables is large, even if it remains smaller than the sample size. They proposed an alternative asymptotic framework. It would be valuable to examine how the methods evaluated here behave under those high-dimensional conditions.
@@
Finally, our comparison addressed model uncertainty arising from uncertainty about which variables to include, but not other forms of model uncertainty such as alternative link functions like the complementary log-log function. Extending the present framework to these settings would provide a more complete picture of model uncertainty in generalized linear models. Likewise, future work could explore the impact of different model space priors. In this study, we used a truncated beta-binomial (1,1) prior that assigns equal probability across model sizes up to n2n-2. This choice of model space prior may partly explain the observed tendency toward larger posterior expected model sizes observed for some BAS methods in low-dimensional datasets. Therefore, alternative priors that more strongly penalize model complexity or encode prior information about sparsity could lead to different conclusions. Finally, it should be noted that the performance differences between BIC.BMA and BIC.BAS partly reflect differences in their model space priors and search algorithms.
Note that, while the datasets in our study are based on real data, the actual observations are simulated from a model in the set of all possible models being considered. Thus the study is in the context of the M-closed framework [9]. One implication is that the generating model is linear in a set of considered predictors. This is more general than one might think, however. For example, nonlinear effects can be included in the framework by including regressors that are nonlinear transformations of regressors or interactions.
@@
In summary, our study provides a comprehensive, preregistered, empirical comparison of 28 established methods for variable selection and inference under model uncertainty in logistic regression. The EB-local prior was the most robust method across separation and non-separation conditions, offering a balance of accuracy, stability, and computational efficiency. Penalized methods, such as LASSO, SCAD and MCP, are strong alternatives when separation is likely and quantifying model uncertainty is deemed unnecessary. Together, these findings and the preregistered, data-based design of the study demonstrate the feasibility of transparent, reproducible simulation studies in statistical methodology.
@@
Method References Package Function
Hyperg-g [72, 91] BAS-V1.7.5 betaprior = hyper.g()
@@
$g=\sqrt{n}$ [30] BAS-V1.7.5 betaprior = g.prior(sqrt(n))
$g=\max(n,\,p^{2})$ [30] BAS-V1.7.5 betaprior = g.prior(max(n, p^2))
$g=4$ [108] BAS-V1.7.5 betaprior = g.prior(4)
CCH [71] BAS-V1.7.5 betaprior = CCH(1, n, 0)
Robust [5] BAS-V1.7.5 betaprior = robust()
Intrinsic [6] BAS-V1.7.5 betaprior = intrinsic()
Beta-prime [75] BAS-V1.7.5 betaprior = beta.prime()
EB-local [22, 40] BAS-V1.7.5 betaprior = EB.local()
EB-global [22, 40] BAS-V1.7.5 betaprior = EB.global()
AIC [2, 14, 71] BAS-V1.7.5 betaprior = aic.prior()
BIC.BAS [87, 71, 85] BAS-V1.7.5 betaprior = bic.prior()
Ridge [53, 67] glmnet-V4.1.9 cv.glmnet(..., alpha = 0)
Elastic net [114] glmnet-V4.1.9 cv.glmnet(..., alpha = 1/2)
LASSO [100] glmnet-V4.1.9 cv.glmnet(..., alpha = 1)
$p$-value [56] stats-V4.5.0 glm(...)
$p<.05$ [56] stats-V4.5.0 glm(...)
$p<.005$ [56] stats-V4.5.0 glm(...)
Forward [56] stats-V4.5.0 step(..., direction = "forward")
Backward [56] stats-V4.5.0 step(..., direction = "backward")
Both [56] stats-V4.5.0 step(..., direction = "both")
Spike and Slab [101, 92] BoomSpikeSlab-V1.2.6 logit.spike(...)
MCP [113, 11] ncvreg-V3.15.0 cv.ncvreg(..., penalty = "MCP")
SCAD [28] ncvreg-V3.15.0 cv.ncvreg(..., penalty = "SCAD")
Induced Smoothed LASSO [20] islasso-V1.5.2 islasso(...)
Firth’s Bias-Reduced [31, 49] logistf-V1.26.1 logistf(...)
BIC.BMA [87] BMA-V3.18.20 bic.glm(...)
Table 1 provides an overview of the included methods. By methods, we also refer to different implementation options within R packages; for instance, the BAS package offers several parameter priors that have been shown to yield varying performance [71, 81], and we therefore treat each of these priors as a distinct method. For each method we provide the generic function call used in our simulations; for Bayesian methods this also indicates the parameter prior distribution used, with the exception of the BMA package, where default prior settings are used. We did not vary the model space prior for any of the Bayesian methods. For all the methods in the BAS package we fixed the model space prior to a truncated beta-binomial($\alpha=1,\,\beta=1$) so that models with n2n-2 or more predictors received a prior probability of 0; for datasets with $n>p$ this reduces to an untruncated beta-binomial prior that is uniform over model size. For the Spike and Slab method we used the default options in the LogitZellnerPrior() function with prior.success.probability = 0.5 and expected.model.size = 1, which assigns each variable a prior inclusion probability of $\min(1,1/p)$. For BIC.BMA we used the default prior inclusion probability of $0.5$ for each variable. In the preregistration, we indicated that we would conduct initial simulation studies to determine which model-space priors to use. However, due to computational considerations, we ultimately decided to retain the default choices described above.
We used the default MCMC and optimization algorithms for all the R packages. For the sake of computational efficiency, we limited the number of posterior samples for BAS and BoomSpikeSlab to 10,000, matching the number of samples used in earlier large scale simulation studies [e.g., 81, 82].
@@
We began by screening online data repositories (e.g., Kaggle, Open Psychometrics, University of California Irvine Machine Learning Repository), datasets available in software packages (e.g., R, the JASP data library 58), educational materials [e.g., 56, 1], and additional resources such as published research articles for suitable empirical datasets. This initial screening stage yielded 51 candidate datasets. From these, we excluded those that were not readily accessible (i.e., required substantial preprocessing or assumptions about the variables) and those that were unsuitable for logistic regression (e.g., highly imbalanced outcome variables or clear assumption violations). This selection process resulted in 24 candidate datasets for the simulation. These candidate data sets correspond to those specified in the preregistration document.
In addition to ensuring diversity in data dimensions and subject areas, we also assessed the suitability of each dataset for the data-generating process (DGP; described in the following subsection), as we anticipated that some might fail to produce a valid data-generating model (DGM). A dataset was automatically included if it (a) produced no computational issues or warnings, (b) exhibited no separation problems (thus requiring no special handling; see the following subsection), (c) yielded a DGM that was a proper subset of the full predictor set, (d) achieved a pseudo-$R^{2}$ greater than $0.1$, and (e) was not too thematically similar to another dataset already included. Based on these criteria, 11 datasets were retained for inclusion, while 8 were excluded for failing one or more conditions.
@@
burn 388 6 [56, 66]
telescope 19,020 10 [38]
dyslexia 118 11 [25, 26]
housing 545 11 [109]
heart 1,025 13 [65]
covid19 468 25 [16]
voting 345 34 [10]
divorce 170 54 [102]
nki70 144 70 [45]
musk 6,598 166 [18]
Singh 102 6,033 [94]
@@
      • ISIS: A method that iteratively selects important variables by ranking them according to their marginal correlations with the outcome [29]. It then performs conditional screening given previously selected variables.
@@
    Simulating outcome variables:
    • After obtaining the coefficients for the selected predictors, we calculate the success probability for each observation in the dataset, that is $\pi_{i}=1/(1+\exp{\left(-\bm{\hat{\beta}}^{\intercal}\mathbf{x}_{i}\right)})$, with $i=1,\ldots,n$, where we bound the linear predictor $\bm{\beta}^{\intercal}\mathbf{x}_{i}$ between 10-10 and $10$ to avoid numerical instability. We then generate a random binary outcome variable for each observation, that is, $y_{i}\sim\text{Bernoulli}(\pi_{i})$, for $i=1,\ldots,n$.
    • Finally, we simulate $m=100$ such binary outcome variables for each empirical dataset (Table 2), while keeping the full empirical predictor matrix $\bm{X}$ fixed.
Analysis & Metrics
@@
. This DGP described in the previous subsection yielded a total of $11\times 100$ simulated datasets. Each dataset was analyzed using all 28 candidate methods (Table 1). To ensure robustness, we applied 5-fold cross-validation when analyzing each simulated dataset with each method.
We monitored whether separation occurred in any dataset at any fold of the cross-validation using the R package detectseparation [63]. Out of the 1,100 simulated datasets, 463 (42%) exhibited separation. For the covid19 dataset, 34 out of 100 simulated datasets showed separation; for heart, 28 out of 100; and for housing, 1 out of 100. All simulated datasets showed separation for the divorce, nki70, voting, and Singh datasets, whereas none exhibited separation for the burn, dyslexia, musk, and telescope datasets.
@@
A complete absence of results (i.e., total failure) for an entire simulation based on a specific dataset for a method occurred only in cases where the data exhibited separation. For the Singh dataset, all replications failed for BIC.BMA, the Induced Smoothed Lasso, Firth’s Bias-Reduced method, and all stepwise and $p$-value–based methods (nine methods in total). For the Voting dataset, all replications failed for BIC.BMA, the Induced Smoothed Lasso, Firth’s Bias-Reduced method, forward selection, and all $p$-value–based methods (seven methods in total). For the divorce and nki70 datasets, the $p<0.05$ and $p<0.005$ methods failed completely (two methods each). Finally, for the covid19 dataset, the $p<0.05$ method failed. Overall, these failures accounted for 7.6% of the entire simulation and 19.2% of all instances involving separation.
The performance measures used to compare the methods are described below. Each measure was averaged across the 100 repetitions for each simulation based on the 11 empirical datasets. We also report two additional metrics that, while not used for ranking, are important for interpreting the results.
In Figs. 1 and 2, all performance metrics were first averaged over repetitions for each dataset–method pair, then standardized as log-ratios relative to a fixed reference method (the Spike–and–Slab). These standardized values were exponentiated for display, such that a score of 1 denotes parity with the reference method, values below 1 indicate better performance, and values above 1 indicate worse performance. The overall Full Score corresponds to the geometric mean across all performance metrics. The Available Score is computed on the subset of metrics provided by each method, and the Partial Score is restricted to the core metrics shared across all methods (RMSE and prediction Brier score). Methods are ranked according to the Partial Score. The additional metrics were not standardized.
@@
  • Mean Interval Score (MIS): The interval score [44] is a measure to assess the quality of confidence/credible intervals for the coefficients. It is calculated as the interval width penalized by the missed coverage when the true value lies outside the interval. The smaller interval score the better.
@@
    where $l$ and $u$ are respectively the lower and upper bound of the confidence/credible interval, and $2/\alpha$ is a penalization coefficient for the missed coverage, with $\alpha=0.05$. We compute the mean interval score (MIS) for all coefficients in a data set.
@@
  • Brier score: The Brier score [12] is a measure used to quantify the accuracy of binary probabilistic predictions on held-out test data. It is defined as the mean squared difference between predicted probabilities and the actual binary outcomes. Lower Brier scores indicate better calibrated and more accurate predictions:
@@
  • CPU time: The time in seconds taken to run an analysis on one dataset using a specific method, (including all cross-validation folds). All methods were run on a single CPU core (with OpenMP and OpenBLAS threading disabled), so no parallelization was used. Methods relying on compiled code internally, still benefit from optimized compiled routines but not from parallelism. As already mentioned, MCMC-based methods (BAS, BoomSpikeSlab) were given a fixed budget of 10,000 iterations; no formal convergence diagnostics were applied, so runtimes for these methods reflect a fixed computational budget rather than a convergence-equivalent comparison with the optimization-based methods. The values presented in Figures 1 and 2 are in minutes. This metric is included as an indicative measure of computational demand; given differences in implementation, algorithmic type, and computational budget, it should not be interpreted as a definitive basis for speed comparison across methods.
  • Failure rate: The proportion of failed models, defined as the proportion of analyses that did not produce any usable output and exited with an error (i.e., complete failure).